There is a supply and demand graph, linear. Equilbirum quantity= 30. Equilibrium price= 10. Quantity is 50 and price is 20. The point on the vertical axis which intercepts witht he supply curve = 5.

At equilbrium, the price elasticity of supply is given by?
1
Expert's answer
2020-10-22T18:13:11-0400
"\frac {Q-50}{5-50}=\frac {p-20}{0-20}"


"Q=2.25 p+5"


"E=Q'_p \times \frac{p}{Q}"

"E=2.25\times\frac{10}{30}=0.75"